OpenStudy (anonymous):

Veronica ordered a decorative garden trellis in the shape of a square pyramid. The online description of the trellis indicated that the base has a side length of 14 inches and height along each triangular face of 30 inches. What is the surface area of Veronica's new garden decoration?

OpenStudy (anonymous):

there is the surface area of the base which is a square of side 14, then 4 triangles with base 14 and height 30. square base: 14^2=196 triangles:4*(1/2)*(14*30)=840 840+196=1036 in^2

OpenStudy (anonymous):

Gracias

OpenStudy (anonymous):

de nada

OpenStudy (anonymous):

Now how did u figure this out?

OpenStudy (anonymous):

Well in the problem description it says we have a square pyramid, which means the base is a square and there are 4 congruent triangles that meet at the tip.

OpenStudy (anonymous):

It gives the side length of the square which is also the base of each triangle, and it gives the slant height or height of each triangle, so we add up all the areas of the parts for total surface area

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

Thnx